OpenStudy (anonymous):

Sally worked the following equation on her paper. She checked her answer and found that is was incorrect. Identify her mistakes and explain to her how to correct them. (hint: There are 2 mistakes.) 3(8x-2)=46-2(12x+1) 24x-6=46-24x-6=48-24x 48x-6=48 48x=42 x=42/48 or 7/8

OpenStudy (anonymous):

3(8x-2)=46-2(12x+1) On the RHS of the above expression -2 times +1 is -2, not -6. 48x-6=48 When -6 is moved to the RHS, the result is 48+6, 54, not 42.